Find an equation of the sphere that has center (2,3,-1) and contains the point (1,7,-9)


1
Expert's answer
2021-12-14T14:05:13-0500

The equation of the sphere is


"(x-2)^2+(y-3)^2+(z+1)^2=R^2"

The sphere contains the point "(1,7,-9)"


"(1-2)^2+(7-3)^2+(-9+1)^2=81"

The equation of the sphere that has center "(2,3,-1)" and contains the point "(1,7,-9)"

"(x-2)^2+(y-3)^2+(z+1)^2=81"